OpenStudy (anonymous):

Which is a solution of the absolute value function y = |x - 4| +2. A. (4, 2)<--- my answer B. (2, 4) C. (-4, 2) D. (2, -4)

OpenStudy (nincompoop):

how? can you show how you solved it?

OpenStudy (anonymous):

do u divide

OpenStudy (jdoe0001):

I'm assuming you mean "the vertex of the expression" ?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

|x - 4| <--- what value for "x" makes that = 0?

OpenStudy (anonymous):

|x-4|+2

OpenStudy (jdoe0001):

yes.... but just focusing on the absolute value expression what value of "x" would make the expression 0?

OpenStudy (anonymous):

do you divide them and get 2

OpenStudy (jdoe0001):

well...hmmm 2?

OpenStudy (jdoe0001):

anyhow... well.. no you don't have to divide them.... you'd just get a value for "x" that makes the absolute expression =0 thus |x - 4| <--- what value for "x" makes that = 0?

OpenStudy (anonymous):

oh okay -4

OpenStudy (anonymous):

so will it be c

OpenStudy (anonymous):

0-4=-4

OpenStudy (jdoe0001):

let's see x = -4 |(-4)-4| = | -8| = 8 well. no dice on -4 so what do you think it'd be?

OpenStudy (anonymous):

well i think this is confusing lol

OpenStudy (jdoe0001):

heehe I didn't mean to say set x = 0 I meant to say a value for "x" that makes the expression in the bars =0 so say for example | x -25| if I set x = 25 then the expression becomes => | (25) - 25| => | 0| = 0 so x = 25 makes | x-25| to 0

OpenStudy (anonymous):

25

OpenStudy (jdoe0001):

for |x - 25| yes, that'd make it 0 how about for | x - 4 | ?

OpenStudy (anonymous):

would it be x=4

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

\(\bf |{\color{red}{ 4}} - 4| +{\color{red}{ 2}} \\ \quad \\ ({\color{red}{ 4,2}})\Leftarrow vertex\)

OpenStudy (anonymous):

so my answer was right it was a

OpenStudy (jdoe0001):

well, yes, but now you know how to obtain the vertex

OpenStudy (anonymous):

well thanks you amazing

OpenStudy (jdoe0001):

yw